OpenStudy (ledah):

Solve for n: 3(-3n + 4) +6n = -12

OpenStudy (anonymous):

-9n+12+6n=-12 -3n=-12-12 -3n=-24 multiplying both sides by a negative sign 3n=24 n=24/3 n=8